OpenStudy (anonymous):

The ratio of square H’s length to square I’s length is 3 : 5. If the length of square H is 18 meters, what is the perimeter of square I?

OpenStudy (anonymous):

@Freelance1 @Somy @11mills1513 please help

OpenStudy (somy):

well first the ratio is 3:5 lets find length of I H is 18 I lets say is X 18---3 X----5 X= 5*18/3 X= 30 so now you have length of I as 30 Perimeter = 4a so 4*30 = your answer

OpenStudy (somy):

i did it from my understanding in chemistry, so cross check with someone in maths just in case :)

OpenStudy (anonymous):

ok thank you so much :D

OpenStudy (anonymous):

it was right XD THANK YOU SO MUCH!!! :D

OpenStudy (somy):

oh that's good :) you are welcome :))